OpenStudy (anonymous):

Use elimination to solve each system of equations. 34. x + y = 4 −2x + 3y = 7

OpenStudy (anonymous):

Multiply first equation with 2 and add the two equations

OpenStudy (anonymous):

2x + 2y = 8 -2x +3y = 7 ......................... 5y = 15 = > y =3 now substitute this in first equation and u will get x value :)

OpenStudy (anonymous):